In an air motor cylinder the compressed air has an internal energyof 450kg/kj at the beginning of the expansion and an internal energy of 220kg/kj after expansion. If the work done by air during the expansion is 120kg/kj,calculate the heat flow to and fro the cylinder and state if the system gain or losses heat
1
Expert's answer
2020-05-18T10:04:53-0400

As per the question,

Internal energy of the compressed air "(U_1)=450kg\/KJ"

Internal energy after expansion"(U_2)=220 kg\/kJ"

Work done by air during expansion"(W)=120 kg\/KJ"

"Q=W+(U_2-U_1)"

"Q=120+(220-450)=120-230=-110kg\/KJ"

Here negative sign is showing that heat rejected by the air is 110 kg/KJ, so here heat is getting loss.